An airplane covered 14 miles of its route while decreasing its altitude by 13,000 feet. Find the slope of the airplane's line of descent. Round to the nearest hundredth. [Hint: 1 mi = 5280 feet.]
Answer by checkley75(3666)

SET THE ORIGINAL POINT @ (0,0)
THE X VALUE WOULD BE 14*5,280=73,920
THE Y VALUE WOULD BE -13,000 OR (73,920,-13,000)
THE SLOPE IS (Y2-Y1)/(X2-X1)
SLOPE=-13,000-0)/(73,920-0)
SLOPE=-13,000/73,920
SLOPE=-.176